OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. (9x-3)-3

OpenStudy (lgbasallote):

hint: \[\huge a^{-m} \implies \frac{1}{a^m}\] so what do you think is the value of \[\huge (9x- 3)^{-3}\]

OpenStudy (anonymous):

1/(9x-3)^3

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

problem solved?

OpenStudy (anonymous):

so its simplified?? thats it thats all???

OpenStudy (lgbasallote):

i believe so...unless you'd like to expand that denominator you ahve

OpenStudy (anonymous):

but guess what?? I just looked at the problem again, it didnt copy over right. it should be (9x^3)^3 does it still apply

OpenStudy (lgbasallote):

well kinda...

OpenStudy (lgbasallote):

is that \[(9x^3)^3\] or \[(9x^3)^{-3}\]

OpenStudy (anonymous):

the x has a negative 3 too

OpenStudy (lgbasallote):

so it's \[\huge (9x^{-3})^{-3}?\]

OpenStudy (anonymous):

yes thats it, so will i still use the same equation you gave at first

OpenStudy (lgbasallote):

yep \[\huge a^{-m} = \frac {1}{a^m}\]

OpenStudy (lgbasallote):

so what do you have?

OpenStudy (anonymous):

yeah i did 1/(9x^3)^3

OpenStudy (lgbasallote):

dont you mean \[\huge \frac{1}{(9x^{-3})^3}\]

OpenStudy (anonymous):

ohhhhh so I am only changing the outer exponent??

OpenStudy (lgbasallote):

so far yes.

OpenStudy (lgbasallote):

now use the rule \[\huge (a^b x^y)^c \implies a^{bc} x^{yc}\]

OpenStudy (anonymous):

wow okay 729x^-27 ?????

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now use \[\huge a^{-m} \implies \frac{1}{a^m}\] on x^-27

OpenStudy (anonymous):

1/x^27

OpenStudy (lgbasallote):

right...so all in all that would be..?

OpenStudy (lgbasallote):

here's a hint: 2x^-3 would be 2/x^3 so what would be 729 x^-27

OpenStudy (anonymous):

729x^27

OpenStudy (lgbasallote):

nope...remember what a^-m is

OpenStudy (lgbasallote):

like i said \[\huge 2x^{-3} \implies \frac{2}{x^3}\] so what's \(729 x^{-27}\)

OpenStudy (anonymous):

1/729x^27